OpenStudy (anonymous):

What is the length of MO? 1 ,√3, 2 , 3

OpenStudy (anonymous):



OpenStudy (phi):

if that is a right triangle with hypotenuse = 2 you can use pythagoras a^2 + b^2 = c^2 in this case c (the hypotenuse) is 2, and one leg (call it leg b) is 1 so you have a^2 + 1^2 = 2^2 or a^2 +1 = 4 can you finish ?